Check my work 3 Required information IThe following information applies to the questions displayed below.) The stockholders equity of TVX Company at the beginning of the day on February 5 follows: Part 2 of 3 points Conmon stock-$10 par value, 150,000 shares authorized, 75,000 shares issued and outstanding Paid-in capital in exceas of par valuc, common stock Retained earnings Total atockholders cguity $ 750,000 525,000 675, 000 $1,950, 000 eBook On February 5, the directors declare a 14% stock dividend distributable on February 28 to the February 15 stockholders of record. The stocks market value is S34 per share on February 5 before the stock dividend. The stocks market value is $30 per share on February 28. Hint Print 2. One stockholder owned 750 shares on February 5 before the dividend. Compute the book value per share and total book value of this stockholders shares immediately before and after the stock dividend of February 5. (Round your Book value per share answers References to 3 decimal places.) Before After Book value per share Total book value of shares
Answer #1

Solution 2:

Nos of shares before stock dividend = 75000 shares

Nos of shares after stock dividend = 75000*114% = 85500 shares

Book value per share before stock dividend = Total stockholder's equity / Nos of shares outstanding = $1,950,000 / 75000 = $26 per share

Book value per share after stock dividend = Total stockholder's equity / Nos of shares outstanding after stock dividend = $1,950,000 / 85500 = $22.807 per share

Particulars Before After
Book value per share $26.00 $22.807
Nos of shares held 750 855
Total book value of shares $19,500 $19,500